# Differential operators, anisotropy, and simplicial spheres

### v1

@@ 3.1. Applications in characteristic p
Let $h\in\overline{H}^{m}(\Delta)$ for some $m\leq d/p$. Let $I$ be a $d\times n$ matrix of nonnegative integers such that all rows sums are $p-1$. Let $J$ be a $1\times n$ matrix of nonnegative integers such that the row sum is $d-pm$. Then
$\partial^{I}\deg(h^{p}\cdot x^{J})=(-1)^{d}\deg(h\cdot(x^{I}\cdot x^{J})^{1/p})^{p},$
where $(x^{I}\cdot x^{J})^{1/p}$ is the $p$th root of $x^{I}\cdot x^{J}$ if $x^{I}\cdot x^{J}$ is a $p$th power, and is zero otherwise.
@@
Write $h=\sum_{L}\lambda_{L}x^{L}$ for some $1\times n$ matrices of nonnegative integers $L$ and some $\lambda_{L}\in K$. Using Theorem 1.1 and that partial derivative are linear over $p$th powers, we compute
$\partial^{I}\deg(h^{p}\cdot x^{J})$ $=\sum_{L}\partial^{I}\lambda_{L}^{p}\deg(x^{pL}\cdot x^{J})$
$=\sum_{L}\lambda_{L}^{p}\partial^{I}\deg(x^{pL}\cdot x^{J})$
$=(-1)^{d}\sum_{L}\lambda_{L}^{p}\deg(x^{L}\cdot(x^{I}\cdot x^{J})^{1/p})^{p}$
$=(-1)^{d}\deg(h\cdot(x^{I}\cdot x^{J})^{1/p})^{p}.\qed$

### v2

@@ 3.1. Applications in characteristic p
Let $h\in\overline{H}^{m}(\Delta)$ for some $m\leq d/p$. Let $J$ be a $1\times n$ matrix of nonnegative integers such that the row sum is $d-pm$. Then
$\deg(h^{p}\cdot x^{J})=(-1)^{d}\sum_{I}\deg(h\cdot(x^{I}\cdot x^{J})^{1/p})^{p}\frac{a^{I}}{I!},$
where $I$ varies over all $d\times n$ matrices of nonnegative integers whose row sums are all $p-1$. Moreover, if $I$ is a $d\times n$ matrix of nonnegative integers such that all rows sums are $p-1$, then
$\partial^{I}\deg(h^{p}\cdot x^{J})=(-1)^{d}\deg(h\cdot(x^{I}\cdot x^{J})^{1/p})^{p}.$
@@
The second identity follows by applying $\partial^{I}$ to the first identity, and using the fact that partial derivatives commute with $p$th powers in characteristic $p$.
We prove the first identity. Write $h=\sum_{L}\lambda_{L}x^{L}$ for some $1\times n$ matrices of nonnegative integers $L$ and some $\lambda_{L}\in K$. Using Theorem 1.1, we compute
$\deg(h^{p}\cdot x^{J})$ $=\sum_{L}\lambda_{L}^{p}\deg(x^{pL}\cdot x^{J})$
$=(-1)^{d}\sum_{L}\lambda_{L}^{p}\sum_{I}\deg(x^{L}\cdot(x^{I}\cdot x^{J})^{1/p})^{p}\frac{a^{I}}{I!}$
$=(-1)^{d}\sum_{I}\deg(h\cdot(x^{I}\cdot x^{J})^{1/p})^{p}\frac{a^{I}}{I!}.\qed$